OpenStudy (anonymous):

John is now 3 times as old as his brother Sam. In 5 years, John will be twice as old as Sam will be then. Find their present ages.

OpenStudy (anonymous):

Solve the following for j and s: { j = 3 s, (j + 5) =2 ( s + 5) } j = 15, s = 5

OpenStudy (lalaly):

fhemteeh wella aktobo ana?

OpenStudy (anonymous):

uktuby if u want

OpenStudy (lalaly):

lol ente bedek?

OpenStudy (anonymous):

yea lol

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (lalaly):

let johns age=j sams age=s so johns age after 5 years=j+5 w nafs el shee la sams age bel nsbeh la John is now 3 times as old as his brother Sam j=3s w la In 5 years, John will be twice as old as Sam will be then j+5=2(s+5) solve that equation bema enno j=3s substitute 3s in j in the second equation 3s+5=2(s+5) 3s+5=2s+10 s=5 so sams age is 5 and johns age is 5*3= 15

OpenStudy (lalaly):

3am be7dar arab idol, eshee kteer ghabee lol

OpenStudy (anonymous):

LOL I NEVER WATCHED IT I WATCHED ARAB GOT TALENT

OpenStudy (lalaly):

be7aznoo lool today awwl 7al2a el youm,

OpenStudy (lalaly):

lol forget abt (elyoum) lol

OpenStudy (anonymous):

u live in the blad?

OpenStudy (lalaly):

no bas 3ndna mbc 3l tv, men el net mashbook 3l tv xD

OpenStudy (anonymous):

yah me too lol

OpenStudy (lalaly):

yeah its on mbc1 :P bas ma bensa7ek sara7a

OpenStudy (anonymous):

wat time does it come on in california?

OpenStudy (lalaly):

ma ba3ref, its on now mo kelhom nafs il wa2t taba3 mbc?

OpenStudy (anonymous):

oh ok i will check